OpenStudy (anonymous):

w/4 + 4x - 1/5 = 1/8 I need help solving this

OpenStudy (anonymous):

Are you solving for x?

OpenStudy (anonymous):

I think so...Not quite sure...let me check to see if it tells me.

OpenStudy (anonymous):

When solving a rational equation the first step that we must always take would be to make sure that you convert the equation to an equation without denominators. This is the explanation I got.

OpenStudy (anonymous):

OK that makes sense, but it still doesn't tell us whether we should solve for x or w. Or are we just supposed to convert the equation and not solve anything? Any more hints in your instructions?

OpenStudy (anonymous):

That is all I got..I get so frustrated. I guess we have to find the problems in this and address them so they know what they are

OpenStudy (anonymous):

Shall we try solving for x then?

OpenStudy (anonymous):

I guess...just so you know...to give a heads up..I am horrible at this stuff...I get so confused

OpenStudy (anonymous):

No problem, I'll try my best to help you understand :)

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

Do you have a guess as to how to start, or should I show the first step?

OpenStudy (anonymous):

This week is a confusing week for me...just as I get the hang of things I get confused then frustrated...

OpenStudy (anonymous):

ok so scratch that question...

OpenStudy (anonymous):

That's ok, it happens to all of us. The first thing we need to do is find the Least Common Denominator (the LCD). Do you know how to do that

OpenStudy (anonymous):

I have one that I did, but he wants the original rational equation of this 10 + 4(x – 2) = 5(x) 10 + 4x – 8 = 5x 4x + 2 = 5x 2 = x or x= 2

OpenStudy (anonymous):

I thought I had did this...and that this answer was it...apparently not....lol

OpenStudy (anonymous):

Hang on, I'm confused. Does this have anything to do with the first problem you posted or have we moved on?

OpenStudy (anonymous):

moved on...I just asked one of the girls in my class by text and she said not to worry about the first one...fix the stuff we did wrong individually...sorry

OpenStudy (anonymous):

OK. So he wants the original rational equation of 10+4(x-2)=5x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I thought I had done this..with all the work I did...not so much he said

OpenStudy (anonymous):

Well, a rational equation is an equation that has fractions in it. I think he wanted you to manipulate the equation he gave to look like one. Instead, what you did was solve for x. More coming...

OpenStudy (anonymous):

ok see now I am getting confused...

OpenStudy (anonymous):

What I think you have to do is distribute the 4 like you did so that you get 10+4x-8=5x, then divide both sides of the equation by 40 to get a rational equation

OpenStudy (anonymous):

Can you maybe tell me more about what you're confused about in specific?

OpenStudy (anonymous):

most of it....I am just not a math person at all....not sure why it never clicked for me.

OpenStudy (anonymous):

I hate the fact I have to ask for help...

OpenStudy (anonymous):

Don't worry about it. Math is chalenging and everyone struggles with it. Plus, it's confusing when you're being tutored online

OpenStudy (anonymous):

Back to the problem...

OpenStudy (anonymous):

sure...I will be checking in a few at a time..I have 2 very sick little girls...sorry if I do not respond right away... I will be back if not:)

OpenStudy (anonymous):

So, a rational equation is an equation that has fractions in it. Like this one (just a random example): x/2+1=x/8

OpenStudy (anonymous):

We want to make the equation you were given look something like that.

OpenStudy (anonymous):

According to your instructions anyway

OpenStudy (anonymous):

So here's what I would do. Take the equation 10 + 4(x – 2) = 5(x)

OpenStudy (anonymous):

Sorry...

OpenStudy (anonymous):

Step 1. Find the least common multiple of 10, 4, and 5. It is 20 because 20 is the smallest number that can divide evenly into 10, 4, and 5

OpenStudy (anonymous):

No problem

OpenStudy (anonymous):

Step 2. Divide every term in the equation by 20. It will look like this: 10/20 + 4(x – 2)/20 = 5(x)/20

OpenStudy (anonymous):

Step 3. Simplify the fractions. It will look like this: 1/2+(x-2)/5=x/4

OpenStudy (anonymous):

And this looks like a rational equation, so I believe this is what your teacher wants

OpenStudy (anonymous):

Thank you so much!! I think I understand this better:) Your great!

OpenStudy (anonymous):

You're welcome. :) So glad I could help. I'm kinda new to this online tutoring thing, so I was nervous.

OpenStudy (anonymous):

Do you have other questions or more questions about this problem?

OpenStudy (anonymous):

no...thank you so much:)

OpenStudy (anonymous):

You're welcome :)